Demand for Corn Flakes is: P = 10 - Q. Supply of Kellogg's Corn Flakes is: P = 2 + Q. Now a generic company enters the market, selling generic Corn Flakes for $3. Assume consumers are indifferent between generic and Kellogg's Corn Flakes.

When the generic corn flakes enter the market, Kellogg's will sell how many less boxes of their own cereal?

Enter as an absolute value (NOT a negative number).
Answer #1

Ans: From the above question. kellogg's equili\brium price and quantity can be calculated by equating the supply and demand equations,

10-Q=2+Q

10-2= 2Q

Q=4

Price= 10-Q = 10-4= 6

Now, Kellogg's quantity = 4 and price = 6

Generic corn flakes sell their product at $3

Substituting this value in the demand equation,

3= 10-Q = Q= 7

Therefore, Kellogg's will be selling 3 fewer boxes in the market i.e. (7-4 =3)